# Learning How to Vote With Principles Axiomatic Insights Into the Collective Decisions of Neural Networks

### v1

@@ Abstract
Can neural networks be applied in voting theory, while satisfying the need for transparency in collective decisions? We propose axiomatic deep voting: a framework to build and evaluate neural networks that aggregate preferences, using the well-established axiomatic method of voting theory. Our findings are: (1) Neural networks, despite being highly accurate, often fail to align with the core axioms of voting rules, revealing a disconnect between mimicking outcomes and reasoning. (2) Training with axiom-specific data does not enhance alignment with those axioms. (3) By solely optimizing axiom satisfaction, neural networks can synthesize new voting rules that often surpass and substantially differ from existing ones. This offers insights for both fields: For AI, important concepts like bias and value-alignment are studied in a mathematically rigorous way; for voting theory, new areas of the space of voting rules are explored.***The source code will eventually be made available here: https://github.com/LevinHornischer/AxiomaticDeepVoting.
Keywords
Computational social choice, machine learning, neural networks, voting theory, axiomatic method, semantic loss function, data augmentation.
1 Introduction
@@
Artificial intelligence (AI) is increasingly applied in many domains, including not just scientific and technological but also societal domains. This poses a dilemma when it comes to social choice, i.e., voting, preference aggregation, and other processes of collective decisions. On the one hand voting systems should be transparent, but the neural networks on which modern AI is built are notoriously opaque. On the other hand neural networks could unearth novel and tailor-made collective decision procedures. Already, state-of-the-art techniques for alignment of Large Language Models (LLMs) with human values—like RLHF (Bai et al.,, 2022) or DPO (Rafailov et al.,, 2024)—rely on the aggregation of human preferences about the generated outputs to fine-tune LLMs. This triggered recent research in guiding such AI alignment using social choice (Conitzer et al.,, 2024).
@@
In this paper, we study how neural networks aggregate votes and preferences. When they form such collective decisions, do they adhere to the normative principles that social choice theory formulated as axioms? This is fundamental both for a discussion of the dilemma and for using social choice for AI alignment. Moreover, it offers new insights for both AI and voting theory. For AI, this provides a rich testing ground to study pressing machine learning concepts like bias, value-alignment and interpretability in a mathematically rigorous way. For example, a network is not biased towards specific individuals if it aggregates their preferences in accordance with the axiom of anonymity; the so-called Pareto principle requires the neural network to align with any preference shared among all individuals; and the well-known axiom of independence entails a certain compositional interpretability of the network. For voting theory, axiomatic deep voting provides a new method for the central quest of exploring the space of voting rules.
Three stylized people utter their preferences (over apples, bananas, and cherries), which are then fed into a neural network, which in turn chooses the apple as a winner. Two stylized scientists ask if this process is anonymous, biased, or cycle-free.
Figure 1: Can neural networks learn to vote with principles?
@@
How are individual preferences best turned into a collective decision? This question is studied by social choice theory (Brandt et al.,, 2016; List,, 2022) and, specifically, voting theory (Zwicker,, 2016). A voting rule is a function that takes as input a profile—i.e., a list of each individual’s preferences among a given set of alternatives—and produces as output a collective decision, i.e., the alternative(s) that the rule takes to be most preferred for the group as a whole (see Section 3 for the formal definitions). The most straightforward rule is Plurality (which picks the alternative that is considered best by the most individuals); other classic rules include Borda and Copeland, while a more recent suggestion is Split Cylce.
@@
To study the collective choices of neural networks, we develop the axiomatic deep voting framework (sketched in Figure 1). Deep neural networks are (parametrized) functions that map vectors (typically of a high dimension) to vectors (typically of a low dimension). So, after suitably encoding profiles and collective decisions as vectors, neural networks realize voting rules, i.e., functions from profiles to collective decisions. Discovering a voting rule can then be seen as an optimization problem: updating the neural network parameters until a given desired property is fulfilled. We evaluate a trained neural network in terms of accuracy and axiom satisfaction. While the former is standard in machine learning, the latter is specific to voting theory and its axiomatic method (Thomson,, 2001; List,, 2011). Different axioms describe different desirable properties of voting rules. An example is the already mentioned anonymity axiom which requires that the names of the voters should not influence the collective decision.
@@
With this framework, we investigate three specific questions.
@@
    Correct for the right reasons? Neural networks can accurately learn standard voting rules, but do they adhere to the normative principles expressed by voting-theoretic axioms?
We observe eminent violations of the axioms, despite high accuracy in mimicking voting rules. So we focus on teaching neural networks the expert knowledge expressed by axioms. There are two common ways to do this. The first is via dataset augmentation (Xia,, 2013):
@@
    Learning principles by example? Can neural networks be trained to adhere to voting-theoretic axioms by training with data exemplifying the axioms?
The second way is via semantic loss functions (Xu et al.,, 2018). For this, we develop a translation of the axioms into loss functions; so, by optimizing this loss during training, the network increases the corresponding axiom satisfaction. Importantly, though, perfect axiom satisfaction is impossible according to the infamous theorem by Arrow, (1951). So we search for the best possible axiom satisfaction:
@@
We test three paradigmatic neural network architectures: multi-layer perceptrons, convolutional neural networks, and word embedding based classifiers. We also check four standard distributions of voter preferences. We find the following:
@@
    Our employed architectures demonstrate similar behavior both regarding accuracy and axiom satisfaction. Importantly, despite high accuracy, they markedly violate critical axioms like anonymity—yet, the news is not as bad for other axioms.
@@ 2 Related Work
2.1 Axiomatic Evaluation of Voting Rules
@@
Overall, our work aligns with the traditional concept of evaluating voting rules based on axioms. However, we also consider learning voting rules and not just evaluating existing ones.
2.2 Neural Networks and Voting
@@
Similarly to all these works, our first experiment considers precisely the problem of using neural networks to learn existing rules from voting theory. However, we systematically study this with axioms: instead of only targeting the right outcomes, we test whether they are obtained via the right principles.
@@
In an initial exploration towards the same direction, Armstrong and Larson, (2019) use a single axiom—prescribing the election of a Condorcet winner when one exists—to train normatively appealing neural networks. This work relies on real data from Canadian federal elections, while ours builds on extensive synthetic data. Additionally, our third experiment illustrates original interactions between sets of different axioms that have not been explored in the literature yet. However, one of the more intricate voting principles not tackled in our paper is fairness. After observing a theoretical trade-off between fairness and certain notions of economic efficiency (some related to the Condorcet winner), Mohsin et al., (2022) train two machine learning models on synthetic data and discover new voting rules that compete well against both Plurality and Borda. Although rule synthesis and axiomatic analysis is not a main focus of that work, the obtained results enforce the idea that machine learning methods can beat existing ones from economic theory when optimized for principled learning.
@@
Other promising lines of research target learning an abstract voting rule given examples about its choices (Procaccia et al.,, 2009) and designing a voting rule that maximizes some notion of social welfare (Anil and Bao,, 2021). Holliday et al., (2024) explore the strategic manipulation of voting rules by MLPs of different sizes, generating elections of up to 6 alternatives and 21 voters. They find that sufficiently large MLPs learn to profitably manipulate all examined voting rules only with information about the pairwise majority victories between alternatives. But some rules like Split Cycle seem more resistant than other rules (e.g., Plurality and Borda).
@@
A different approach, rather orthogonal to ours, is to consider AI models as the individuals who vote, instead of using them as the aggregation mechanisms. In this vein, Yang et al., (2024) consider a human voting experiment with 180 participants to establish a baseline for human preferences and conducted a corresponding experiment with LLM (e.g., GPT-4) agents. The voting behavior of the networks seems to be affected by the presentation order of the alternatives, as well as the numerical ID assigned to each LLM representing a voter. Some voting rules such as Borda show that LLMs may lead to less diverse collective outcomes. Importantly GPT-4 seems to over-rely on stereotypical demographics of the voters it is supposed to mimic. Similarly, using data from Brazil’s 2022 presidential election, Gudiño Rosero et al., (2024) tests the accuracy with which LLMs predict an individual’s vote. They find that LLMs are more accurate than a naive rule guessing that individuals simply vote for the proposals of the candidate most aligned with their political orientation.
2.3 Social Choice for AI Alignment
@@
A growing research area studies how social choice theory can be used to guide the alignment of modern AI methods with human values and moral judgments. Conitzer et al., (2024) highlight a series of technical connections—for example, the alternatives in a voting context could be treated as all possible parameterizations of a network, or as all its possible answers. As an indication, in a popular work about a controversial topic, Noothigattu et al., (2018) use data from the online ‘moral machine experiment’ to build a model of aggregated moral preferences aimed at guiding the decision making of autonomous vehicles. On a more theoretical level, Mishra, (2023) utilize Arrow’s theorem to prove that there does not exist any AI system that can treat all its users and human supervisors equally. We do not directly engage with the ethical dimension of this research area; still, we participate in the related foundational discussion by studying whether neural networks can learn to vote with principles.
3 Preliminaries on Voting Theory
@@
We work in the standard setting of voting theory, where a finite set $N$ of voters $N$ have preferences, which are linear orders (also called rankings) over a finite set $A$ of alternatives (Zwicker,, 2016). Set $m:=|A|$ and $n:=|N|$. We denote by $\boldsymbol{P}=(P_{1},...,P_{n})$ a preference profile, i.e., a vector with the preference $P_{i}$ for every voter $i\in N$. This is illustrated in Figure 2.
@@
Figure 2: A voting profile, with voters $N=\{1,\ldots,6\}$ and alternatives $A=\{a,b,c,d\}$. Each column depicts the preference of the individual voter; e.g., voter 1 prefers alternative $c$ most, followed by alternative $a$, etc.
@@
For a permutation of the alternatives $\sigma:A\to A$, the ranking $\sigma(P)$ is obtained by applying $\sigma$ elementwise to the ranking $P$, and $\sigma(\boldsymbol{P})=(\sigma(P_{1}),\ldots,\sigma(P_{n}))$. For a permutation of the voters $\pi:N\to N$, we define $\pi(\boldsymbol{P})=(P_{\pi(1)},...,P_{\pi(n)})$. A voting rule is a function $F$ that determines the winning alternatives for each such profile. Formally, $F:\boldsymbol{P}\mapsto S$, where $\emptyset\neq S\subseteq A$ [^1] We use the Python package pref-voting in all our experiments.
3.1 Voting Rules
@@
Voting rules usually fit into one of two categories: scoring rules and tournament solutions. Scoring rules assign a score to each alternative depending on its position in the linear preference of each voter and declare as winners those alternatives with the highest score across all voters. The two primary scoring rules are Plurality (assigning score 1 to an alternative each time it is ranked first by a voter, and score 0 otherwise) and Borda (assigning score $m-1$ to an alternative ranked first by a voter, score $m-2$ to an alternative ranked second, and so on, until score 0 is assigned to an alternative ranked last by a voter). [^2] Plurality and Borda are often contrasted in voting (Hatzivelkos,, 2018; Terzopoulou,, 2023).
@@
Tournament solutions on the other hand are based on tournaments that capture pairwise comparisons between the alternatives, induced by the voters’ preferences. For $x,y\in A$, let $N^{\boldsymbol{P}}_{x\succ y}$ be the set of voters $i$ in the profile $\boldsymbol{P}$ that consider $x$ better than $y$ in $P_{i}$, and $n^{\boldsymbol{P}}_{x\succ y}:=|N^{\boldsymbol{P}}_{x\succ y}|$. A characteristic tournament solution is the Copeland rule, which selects as winners the alternatives that beat the most other alternatives in a pairwise majority contest: $argmax_{x\in A}|\{y\in A:n^{\boldsymbol{P}}_{x\succ y}\geq n^{\boldsymbol{P}}_{y\succ x}\}|$. The weighted tournament of a profile is a weighted directed graph the nodes of which are alternatives with an edge from $x$ to $y$ of weight $n^{\boldsymbol{P}}_{x\succ y}$. Suppose that in each cycle of the graph, we simultaneously delete the edges with minimal weight. Then the alternatives with no incoming edges are the winners of Split Cycle (Holliday and Pacuit, 2023a, ). If there is only one Split Cycle winner in a profile $\boldsymbol{P}$, then this also is the winner of Stable Voting; otherwise $x$ is a winner of Stable Voting if for some alternative $y$ it holds that $x$ is a Split Cycle winner with the maximal margin $n^{\boldsymbol{P}}_{x\succ y}$ such that $x$ is a Stable Voting winner in the profile $\boldsymbol{P}_{-y}$ obtained from $\boldsymbol{P}$ after deleting alternative $y$ (Holliday and Pacuit, 2023b, ).
@@
Two prominent rules that do not fit into the two above categories are Blacks and Weak Nanson. Black returns the Condorcet winner (i.e., the alternative beating every other alternative in a pairwise strict majority contest) if one exists, otherwise it returns the Borda winners. Weak Nanson is defined iteratively on voting profiles of various sizes. In each round, all alternatives with Borda score at most as high as the average Borda score are removed. Whenever all alternatives have the same Borda score, they all win; otherwise the alternative that remains in the last round wins. In our third experiment, we use even more common voting rules for detailed comparisons, though they are not essential to the paper; hence we refer to the introductory chapter of Zwicker, (2016) and to the pref-voting documentation for the relevant definitions.
3.2 Axioms

### v2

# Learning How to Vote with Principles: Axiomatic Insights Into the Collective Decisions of Neural Networks

@@ Abstract.
Can neural networks be applied in voting theory, while satisfying the need for transparency in collective decisions? We propose axiomatic deep voting: a framework to build and evaluate neural networks that aggregate preferences, using the well-established axiomatic method of voting theory. Our findings are: (1) Neural networks, despite being highly accurate, often fail to align with the core axioms of voting rules, revealing a disconnect between mimicking outcomes and reasoning. (2) Training with axiom-specific data does not enhance alignment with those axioms. (3) By solely optimizing axiom satisfaction, neural networks can synthesize new voting rules that often surpass and substantially differ from existing ones. This offers insights for both fields: For AI, important concepts like bias and value-alignment are studied in a mathematically rigorous way; for voting theory, new areas of the space of voting rules are explored.
doi: 10.1613/jair.1.18890journalvolume: 83article: 25publicationmonth: 8journalyear: 2025
1. Introduction
@@
Artificial intelligence (AI) is increasingly applied in many domains, including not just scientific and technological but also societal problems. This poses a dilemma when it comes to social choice, i.e., voting, preference aggregation, and other processes of collective decisions. On the one hand, voting systems should be transparent, but the neural networks on which modern AI is built are notoriously opaque. On the other hand, neural networks could unearth novel and tailor-made collective decision procedures. Already, state-of-the-art techniques for alignment of Large Language Models (LLMs) with human values—like RLHF [^1] Bai et al., 2022. Training a Helpful and Harmless Assistant with Reinforcement Learning from Human Feedback. arXiv:2204.05862. or DPO (Rafailov et al., 2024)—rely on the aggregation of human preferences about the generated outputs to fine-tune LLMs. This triggered recent research in guiding such AI alignment using social choice (Conitzer et al., 2024).
@@
In this paper, we study how neural networks aggregate votes and preferences. When they form such collective decisions, do they adhere to the normative principles that social choice theory formulates as axioms? This is fundamental both for a discussion of the dilemma and for using social choice for AI alignment. Moreover, it offers new insights for both AI and voting theory. For AI, this provides a rich testing ground to study pressing machine learning concepts like bias, value-alignment and interpretability in a mathematically rigorous way. For example, a network is not biased towards specific individuals if it aggregates their preferences in accordance with the axiom of anonymity; the so-called Pareto principle requires the neural network to align with any preference shared among all individuals; and the well-known axiom of independence entails a certain compositional interpretability of the network. For voting theory, axiomatic deep voting provides a new method for the central quest of exploring the space of voting rules.
[…]
A stylized neural network with preferences of three agents inputted on the left and a winning set outputted on the right.
Figure 1. Can neural networks learn to vote with principles?
@@
How are individual preferences best turned into a collective decision? This question is studied by social choice theory (Brandt et al., 2016; List, 2022), and specifically voting theory (Zwicker, 2016). A voting rule is a function that takes as input a profile—i.e., a list of each individual’s preferences among a set of alternatives—and produces as output a collective decision, i.e., the alternative(s) that the rule takes to be most preferred for the group as a whole (see Section 3 for the formal definitions). The most straightforward rule is Plurality (picking the top alternative of most individuals); other classic rules include Borda and Copeland, while a recent suggestion is Stable Voting.
@@
To study the collective choices of neural networks, we develop the axiomatic deep voting framework (sketched in Figure 1). [^2] The source code is available here: https://github.com/LevinHornischer/AxiomaticDeepVoting. Deep neural networks are (parametrized) functions that map vectors (typically of a high dimension) to vectors (typically of a low dimension). So, after suitably encoding profiles and collective decisions as vectors, neural networks realize voting rules, i.e., functions from profiles to collective decisions. Discovering a voting rule can then be seen as an optimization problem: updating the neural network parameters until a given desired property is fulfilled. We evaluate a trained neural network in terms of accuracy and axiom satisfaction. While the former is standard in machine learning, the latter is specific to voting theory and its axiomatic method (Thomson, 2001; List, 2011). Different axioms describe different desirable properties of voting rules. An example is the already mentioned anonymity axiom which requires that the names of the voters should not influence the collective decision.
@@
With this framework, we investigate the general question of how neural networks aggregate preferences via three more specific questions.
@@
    Correct for the right reasons? Neural networks can accurately learn standard voting rules, but do they adhere to the normative principles expressed by voting axioms?
We observe eminent violations of the axioms, despite high accuracy in mimicking voting rules. So we focus on teaching neural networks the expert knowledge expressed by axioms. There are two common ways to do this. The first is via dataset augmentation (Xia, 2013):
@@
    Learning principles by example? Can neural networks be trained to adhere to voting axioms by training with data exemplifying the axioms?
The second way is via semantic loss functions (Xu et al., 2018). For this, we develop a translation of the axioms into loss functions; so, by optimizing this loss during training, the network increases the corresponding axiom satisfaction. Importantly, though, perfect axiom satisfaction is impossible according to the infamous theorem by Arrow (1951). So we search for the best possible axiom satisfaction:
@@
In three experiments, we answer these questions in turn. In each, we test three paradigmatic neural network architectures: multi-layer perceptrons, convolutional neural networks, and word embedding based classifiers. We also check a variety of standard distributions of voter preferences. Our three experiments find, respectively:
@@
    The employed architectures demonstrate similar behavior both regarding accuracy and axiom satisfaction. Importantly, despite high accuracy, they markedly violate critical axioms like anonymity—yet, the news is not as bad for other axioms.
@@ 2. Related Work
2.1. Axiomatic Evaluation of Voting Rules
@@
Overall, our work aligns with the traditional concept of evaluating voting rules based on axioms. However, we also consider learning voting rules and not just evaluating them.
2.2. Neural Networks and Voting
@@
Similarly to all these works, our first experiment considers precisely the problem of using neural networks to learn existing rules from voting theory. However, we systematically study this with axioms: instead of only targeting the right outcomes, we test whether they are obtained via the right principles. In an initial exploration towards the same direction, Armstrong and Larson (2019) use a single axiom—prescribing the election of a Condorcet winner when one exists—to train normatively appealing neural networks. This work relies on real data from Canadian federal elections, while ours builds on extensive synthetic data. Additionally, our third experiment illustrates original interactions between sets of different axioms that have not been explored in the literature yet.
After observing a theoretical trade-off between fairness (in particular anonymity variations, demanding that different types of voters be treated equally) and certain notions of economic efficiency (some related to the Condorcet winner), Mohsin et al. (2022) train two machine learning models on synthetic data and discover new voting rules that compete well against both Plurality and Borda. Although rule synthesis and axiomatic analysis is not a main focus of that work, the obtained results enforce the idea that machine learning methods can beat existing ones from economic theory when optimized for principled learning. In the slightly different framework of probabilistic voting, MLPs are used to learn voting rules that output distributions over the set of alternatives (rather than the certain winning alternatives), for elections of up to 7 alternatives and 29 voters. [^3] Matone et al., 2024. DeepVoting: Learning Voting Rules with Tailored Embeddings. arXiv:2408.13630. It is shown that the discovered rules can lead to novel ones with improved axiomatic properties, after the appropriate embedding of the input and some adjustments of the output. Another work studies the setting of participatory budgeting (PB), where citizens vote about the projects that they would like to implement in their neighbourhoods, while taking into account their respective costs. [^4] Fairstein, Vilenchik, and Gal, 2024. Learning Aggregation Rules in Participatory Budgeting: A Data-Driven Approach. arXiv:2412.01864. Set Transformer neural networks trained on PB instances with both real and synthetic data are found to both learn existing voting rules and to discover new ones that satisfy societal objectives (such as fair representation). Overall, our paper adds to this literature by offering a concrete framework to study combinations of basic axioms from the ML perspective, in the most standard voting setting.
@@
Other promising lines of research target learning an abstract voting rule given examples about its choices (Procaccia et al., 2009) and designing a voting rule that maximizes some notion of social welfare (Anil and Bao, 2021). Holliday et al. (2025) explore the strategic manipulation of voting rules by MLPs of different sizes, generating elections of up to 6 alternatives and 21 voters. They find that sufficiently large MLPs learn to profitably manipulate all examined voting rules only with information about the pairwise majority victories between alternatives. But some rules like Split Cycle seem more resistant than other rules (e.g., Plurality and Borda). Moreover, a systematic study of the connections between social choice and RLHF is conducted by Dai and Fleisig (2024), including fundamental axiomatic notions such as the Condorcet winner.
@@
A different approach, rather orthogonal to ours, is to consider AI models as the individuals who vote, instead of using them as the aggregation mechanisms. In this vein, Yang et al. (2024) consider a human voting experiment with 180 participants to establish a baseline for human preferences and conducted a corresponding experiment with LLM (e.g., GPT-4) agents. The voting behavior of the networks seems to be affected by the presentation order of the alternatives, as well as the numerical ID assigned to each LLM representing a voter. Some voting rules such as Borda show that LLMs may lead to less diverse collective outcomes. Importantly GPT-4 seems to over-rely on stereotypical demographics of the voters it is supposed to mimic. Similarly, using data from Brazil’s 2022 presidential election, Gudiño Rosero et al. (2024) tests the accuracy with which LLMs predict an individual’s vote. They find that LLMs are more accurate than a naive rule guessing that individuals simply vote for the proposals of the candidate most aligned with their political orientation.
2.3. Social Choice for AI Alignment
@@
A growing research area studies how social choice theory can be used to guide the alignment of modern AI methods with human values and moral judgments. Conitzer et al. (2024) highlight a series of technical connections—for example, the alternatives in a voting context could be treated as all possible parameterizations of a network, or as all its possible answers. As an indication, in a popular work about a controversial topic, Noothigattu et al. (2018) use data from the online ‘moral machine experiment’ to build a model of aggregated moral preferences aimed at guiding the decision making of autonomous vehicles. On a more theoretical level, Arrow’s theorem can be utilized to prove that there does not exist any AI system that can treat all its users and human supervisors equally. [^5] Mishra, 2023. AI Alignment and Social Choice: Fundamental Limitations and Policy Implications. arXiv:2310.16048. Based on an investment game where participants are asked to manage wealth, Koster et al. (2022) find that by optimizing for human preferences, a reinforcement learning model can propose a wealth redistribution mechanism that is supported by the majority of participants. We do not directly engage with the ethical dimension of this research area; still, we participate in the related foundational discussion by studying whether neural networks can learn to vote with principles.
3. Preliminaries on Voting Theory
@@
We work in the standard setting of voting theory, where a finite set $N$ of voters have preferences that are linear orders (also called rankings) over a finite set $A$ of alternatives (Zwicker, 2016). Set $m:=|A|$ and $n:=|N|$. We denote by $\boldsymbol{P}=(P_{1},...,P_{n})$ a preference profile, i.e., a vector with the preference $P_{i}$ for every voter $i\in N$. This is illustrated in Figure 2.
@@
Figure 2. A voting profile, with voters $N=\{1,\ldots,6\}$ and alternatives $A=\{a,b,c,d\}$. Each column depicts the preference of the individual voter; e.g., voter 1 prefers alternative $c$ most, followed by alternative $a$, etc.
@@
For a permutation of the alternatives $\sigma:A\to A$, the ranking $\sigma(P)$ is obtained by applying $\sigma$ elementwise to the ranking $P$, and $\sigma(\boldsymbol{P})=(\sigma(P_{1}),\ldots,\sigma(P_{n}))$. For a permutation of the voters $\pi:N\to N$, we define $\pi(\boldsymbol{P})=(P_{\pi(1)},...,P_{\pi(n)})$. A voting rule is a function $F$ that determines the winning alternatives for each such profile. Formally, $F:\boldsymbol{P}\mapsto S$, where $\emptyset\neq S\subseteq A$. [^6] We use the Python package pref-voting in all our experiments (Holliday and Pacuit, 2025).
3.1. Voting Rules
@@
Voting rules usually fit into one of two categories: scoring rules and tournament solutions. Scoring rules assign a score to each alternative depending on its position in the linear preference of each voter and declare as winners those alternatives with the highest score across all voters. The two primary scoring rules are Plurality (assigning score 1 to an alternative each time it is ranked first by a voter, and score 0 otherwise) and Borda (assigning score m1m-1 to an alternative ranked first by a voter, score m2m-2 to an alternative ranked second, and so on, until score 0 is assigned to an alternative ranked last by a voter). [^7] Plurality and Borda are often contrasted in voting (Hatzivelkos, 2018; Terzopoulou, 2023). Another, less popular scoring rule, is Anti-Plurality, which assigns score 0 to an alternative each time it is ranked last by a voter, and score 1 otherwise.
@@
Tournament solutions on the other hand are based on tournaments that capture pairwise comparisons between the alternatives, induced by the voters’ preferences. For $x,y\in A$, let $N^{\boldsymbol{P}}_{x\succ y}$ be the set of voters $i$ in the profile $\boldsymbol{P}$ that consider $x$ better than $y$ in $P_{i}$, and $n^{\boldsymbol{P}}_{x\succ y}:=|N^{\boldsymbol{P}}_{x\succ y}|$. A classical tournament solution is the Copeland rule, which selects as winners the alternatives that beat the most other alternatives in a pairwise majority contest: $\mathrm{argmax}_{x\in A}|\{y\in A:n^{\boldsymbol{P}}_{x\succ y}\geq n^{\boldsymbol{P}}_{y\succ x}\}|$. In other words, an alternative can be thought to be assigned a Copeland score of 1 for every other alternative to which it is majority preferred, with the winner being the alternative with the highest score overall. Analogously, the Llull rule assigns to an alternative a score 1 for every other alternative to which it is majority preferred, and score $1/2$ for every other alternative to which it is majority tied—if there are no majority ties, then the Llull winners coincide with the Copeland winners. The Top Cycle rule selects the smallest set of alternatives such that each alternative in the set is majority preferred to each alternative outside the set. In a sense, this set captures the notion of a Condorcet winner when a single such alternative does not exist. For the Banks rule, we say that a chain $(x,y,z,\ldots)$ in a tournament is a subset of alternatives that are linearly connected by the majority relation (i.e., $x$ is majority preferred to $y$, $y$ is majority preferred to $z$, and so on). Then $x$ is a Banks winner if it is the maximum element of a maximal chain. To define the recently proposed Stable Voting rule (Holliday and Pacuit, 2023b), we first need to describe—the more computationally expensive, and thus left aside in our analysis—Split Cycle (Holliday and Pacuit, 2023a). The weighted tournament of a profile is a weighted directed graph the nodes of which are alternatives with an edge from $x$ to $y$ of weight $n^{\boldsymbol{P}}_{x\succ y}$. Suppose that in each cycle of the graph, we simultaneously delete the edges with minimal weight. Then the alternatives with no incoming edges are the winners of Split Cycle. If there is only one Split Cycle winner in a profile $\boldsymbol{P}$, then this also is the winner of Stable Voting; otherwise $x$ is a winner of Stable Voting if for some alternative $y$ it holds that $x$ is a Split Cycle winner with the maximal margin $n^{\boldsymbol{P}}_{x\succ y}$ such that $x$ is a Stable Voting winner in the profile $\boldsymbol{P}_{-y}$ obtained from $\boldsymbol{P}$ after deleting alternative $y$.
@@
Other prominent rules that do not fit into the two above categories are Blacks, Baldwin, and Weak Nanson. Black returns the Condorcet winner (i.e., the alternative beating every other alternative in a pairwise strict majority contest) if one exists, otherwise it returns the Borda winners. Weak Nanson (resp., Baldwin) is defined iteratively on voting profiles of various sizes. In each round, all alternatives with below-average (resp., the lowest) Borda score are removed. Whenever all alternatives have the same Borda score, they all win; otherwise the alternative that remains in the last round wins. More voting rules can be defined in iterative terms, based on the idea that less popular alternatives in one round be dropped from all preferences in the next round, until some surviving alternative achieves majority support: Plurality with Runoff consists of at most two rounds—if no alternative is ranked on top by a majority of voters in the first round, then the second round selects the majority winner between the two alternatives that achieved the highest Plurality score in the first round; Instant Runoff (resp., Coombs) is such that in each round the alternatives with the lowest Plurality score (resp., Anti-Plurality score) are eliminated.
A rule different in spirit, the Uncovered Set, relies on the idea of undefeated alternatives. Let us say that alternative $x$ left-covers $y$ if all alternatives $z$ that are majority preferred to $x$ are also majority preferred to $y$. Then, $x$ defeats $y$ if $x$ is majority preferred to $y$ and also left-covers $y$. The winners of Uncovered Set are the undefeated alternatives. Finally, the Kemeny-Young rule is based on a notion of distance between the preferences of the voters: it constructs the linear order that minimises the sum of Kendal Tau distances for all voters’ preferences and elects as winner the alternative on the top of that linear order.
Note that all the aforementioned rules are included in our last experiment, which aims at comparing a wide pool of different voting rules that vary in nature. Our other two experiments are focused on the three most standard voting rules, Plurality, Borda, and Copeland, which are the most frequently studied in the literature to date at the intersection of voting and machine learning. For more detailed discussions of the rules, we refer to the introductory chapter of (Zwicker, 2016) and to the pref-voting documentation.
3.2. Axioms